OpenStudy (anonymous):

x² + y² = 5 and dx/dt = -3, then what is dy/dt when x = 1 and y = 2? Please show work

OpenStudy (anonymous):

2xx'+2yy'=0; 2(1)(-3)+2(2)y'=0; -6+4y'=0; y'=6/4=3/2